# TARS [Uncaptioned image]: MinMax Token-Adaptive Preference Strategy for Hallucination Reduction in MLLMs

### v2

@@ 3.4 Minimization Objective in TARS
Intern-VL2.5-7B (Chen et al. 2024e)§ 7.9 54.7 37.1 3.2 3.54 0.26 - - 36.0 9.1
Qwen-VL2.5-8B (Bai et al. 2025)§ 4.6 54.6 21.1 1.3 3.29 0.27 - - 40.7 8.6
DeepSeek-VL2-27B (Wu et al. 2024)§ 2.4 56.6 16.3 0.9 2.84 0.27 - - 10.0 7.0
GPT-4o (Hurst et al. 2024)§ 2.5 60.9 17.6 0.8 3.87 0.24 - - 29.3 6.7
LLaVA-v1.5-7B (Liu et al. 2023b)§ 7.6 51.7 35.4 4.2 2.02 0.61 80.0 61.8 54.0 15.8
+ RLHF (Sun et al. 2024) 8.3 52.2 41.8 4.5 1.93 0.67 82.0 69.3 56.0 15.2
+ RLAIF (Yu et al. 2025) 3.0 50.3 16.5 1.0 2.89 0.42 88.1 88.0 13.7 4.2
+ HALVA (Sarkar et al. 2025a) 6.9 52.8 33.2 3.5 2.12 0.59 87.5 79.6 47.3 14.6
+ DPO (Li, Lin, and Pei 2024) 4.9 56.6 26.4 2.5 2.19 0.61 87.8 82.0 14.0 5.0
+ CHiP-DPO (Fu et al. 2025) 2.9 57.3 19.9 1.0 2.32 0.57 81.1 91.8 7.3 4.3
+ OPA-DPO (Yang et al. 2025) 2.7 47.4 12.5 0.9 2.78 0.46 87.4 86.2 13.3 4.5
@@
LLaVA-v1.5-13B (Liu et al. 2023b)§ 6.7 52.1 32.5 3.5 2.39 0.53 74.6 55.2 50.0 14.5
+ RLHF (Sun et al. 2024) 7.1 51.4 36.3 3.6 2.10 0.67 83.6 71.2 46.7 11.6
+ HALVA (Sarkar et al. 2025a) 6.5 53.4 30.1 3.3 2.28 0.56 86.8 75.6 42.7 12.1
+ DPO (Li, Lin, and Pei 2024) 4.1 56.7 24.3 2.2 2.48 0.50 85.2 84.3 19.0 7.2
+ CHiP-DPO (Fu et al. 2025) 3.8 58.6 20.8 1.7 2.70 0.46 86.6 74.9 30.0 6.2
+ OPA-DPO (Yang et al. 2025) 2.8 48.4 13.5 1.0 3.02 0.40 87.2 80.7 18.3 5.1
@@
4 Experiments
4.1 Experiment Details
Experiment Setups. We evaluate our method using the multimodal large language model LLaVA-v1.5 (Liu et al. 2023b) at both 7B and 13B parameter scales. All evaluations are conducted using greedy decoding with a temperature of 0 to ensure deterministic outputs and reproducibility.
To enable fair comparison, we align our training configuration with the most data-efficient preference optimization baselines. Specifically, we randomly sample 4.8k instances from the RLHF-V-Dataset (Yu et al. 2024), consistent with OPA-DPO (Yang et al. 2025), and adopt the same training strategy as CHiP-DPO (Fu et al. 2025). All models are trained on eight NVIDIA A100 (80GB) GPUs. We set $\alpha=1$ in Equation 5 and $\beta=1$ in Equation 11 for preference optimization. We implement $\varphi(\cdot)$ using both token masking (Mask) and replacement (Replace) strategies in Equation 8, and set the perturbation constraint strength to $\omega=0.1$ in the adversarial min-max formulation Equation 9. We use a frequency-domain loss weight of $\lambda=0.1$ in Equation 12. Full implementation details and ablation studies are provided in the Appendix.
Evaluation Benchmark. We evaluate TARS across both generative and discriminative benchmarks to ensure that hallucination mitigation does not come at the cost of factual grounding. Our evaluation framework includes: AMBER (Wang et al. 2023), MMHal (Sun et al. 2023), OBJHal (Yu et al. 2024), POPE (Li et al. 2023).
Baseline Methods. We compare against two categories:
(1) Advanced multimodal foundation models: Intern-VL2.5-7B (Chen et al. 2024e), Qwen-VL2.5-8B (Bai et al. 2025), DeepSeek-VL2-27B (Wu et al. 2024), and GPT-4o (Hurst et al. 2024) serve as reference benchmarks.
(2) LLaVA-v1.5 with RL techniques: We evaluate multiple RL-based approaches applied to both the 7B and 13B variants of LLaVA-v1.5, including RLHF (Sun et al. 2024), RLAIF (Yu et al. 2025), HALVA (Sarkar et al. 2025a), as well as three state-of-the-art methods based on direct preference optimization (DPO): DPO (Pi et al. 2024), CHiP-DPO (Fu et al. 2025), and OPA-DPO (Yang et al. 2025). A comparison of algorithmic properties is provided in Table 3.
@@ 4.2 Evaluation on Hallucination Benchmarks
Table 2: Ablation results of token-level perturbation (TP), cross-modal alignment score (CAS), and spectral preference alignment (SPA) on the AMBER and OBJHal using 7B scale.
@@
Figure 4: AMBER hallucination rate vs. preference data scale across 7B and 13B models.
@@ 4.3 Ablation Analyses on Component
(1) Token-level perturbation (TP) in Equation 6, which introduces distributional shifts and proves essential for revealing token-level vulnerabilities—its removal increases Cog from 0.4 to 2.5.
@@
LLaVA-v1.5 (Liu et al. 2023b) - -
+ RLHF (Sun et al. 2024) 122k self-reward
+ RLAIF (Yu et al. 2025) 16k LLaVA-Next
+ HALVA (Sarkar et al. 2025a) 22k GPT-4V
+ DPO (Li, Lin, and Pei 2024) 5k self-reward
+ CHiP-DPO (Fu et al. 2025) 5k self-reward
+ OPA-DPO (Yang et al. 2025) 4.8k GPT-4V
@@ 5 Related Work
Multimodal large language models (MLLMs) extend LLMs by integrating visual inputs to support multimodal reasoning (Chen et al. 2024b; Feng et al. 2025; Jain, Yang, and Shi 2024). Typically, visual features are extracted by a vision encoder, aligned through a connector, and processed by the LLM (Liu et al. 2023b; Parekh et al. 2024). Despite strong performance, MLLMs often produce factually incorrect or visually ungrounded outputs, undermining reliability (Bai et al. 2024; Chen et al. 2025). This issue is more severe than in unimodal LLMs (Chen et al. 2024c; Jiang et al. 2024a), mainly due to modality imbalance (Ma et al. 2024; He et al. 2024) and ineffective fusion (Bellagente et al. 2023; Ji et al. 2023). Recent studies attribute these failures to persistent misalignment between multimodal representations and human expectations, rather than model capacity (Chen et al. 2024d; Liu et al. 2024b; Ruan et al. 2025).
A key bottleneck in addressing MLLM hallucinations lies in aligning model outputs with human preferences for factual consistency. Unlike knowledge-intensive pretraining (Chang et al. 2024; McKinzie et al. 2024) and instruction tuning (Chen et al. 2024a; Liu et al. 2023b), recent methods typically leverage small-scale human preference data refined via reinforcement learning (Yu et al. 2024; Casper et al. 2023). Direct preference optimization (DPO)(Rafailov et al. 2023; Pi et al. 2024) has become a leading approach due to its simplicity and effectiveness, demonstrated in ChiP-DPO(Fu et al. 2025) and OPA-DPO (Yang et al. 2025). However, DPO’s reliance on limited data can cause overfitting to superficial linguistic cues (Setlur et al. 2024; Fu et al. 2025), leading to distributional rigidity and reduced adaptability to modality-specific semantics (Ouali et al. 2024; Song et al. 2024). These limitations call for more adaptive alignment strategies that capture token-level variability and cross-modal dependencies.
@@ 6 Conclusion
  • Section 7 details model configurations, training settings for DPO and TARS, and token perturbation procedures.
  • Section 8 presents the min-max optimization algorithm of TARS in pseudocode form.
  • Section 9 includes extended ablations analyzing the perturbation magnitude and spectral regularization strength.
  • Section 10 reports additional benchmark results and fine-grained hallucination metrics.
  • Section 11 discusses model behavior, including sensitivity and design insights.
  • Section 12 showcases qualitative comparisons on representative examples.
7 Implementation Details
7.1 Base Model Setups
We evaluate our method on LLaVA-v1.5 (Liu et al. 2023b) models with 7B and 13B parameters. LLaVA-v1.5 adopts Vicuna-7B/13B (Chiang et al. 2023) as the language backbone and CLIP-ViT-L/14  (Radford et al. 2021) as the vision encoder. The vision encoder also serves as the similarity function $\mathcal{G}(\cdot)$ used in Eq. (9) to compute alignment between visual inputs and text tokens. All experiments are conducted using greedy decoding with a temperature of 0 to ensure deterministic outputs and reproducibility.
7.2 DPO Training Setups
For fair comparison, DPO (Wang et al. 2024a), CHiP (Fu et al. 2025), and TARS follow the same training protocol as described in CHiP (Fu et al. 2025). Specifically, we set the number of epochs to 3, learning rate to 5e-7, warmup ratio to 0.03, maximum sequence length to 2048, and gradient clipping threshold to 20.0. Notably, TARS requires no task-specific hyperparameter tuning and demonstrates generalization across different base models and datasets. All experiments are conducted on 8$\times$×A100 GPUs (80GB). Each training run takes approximately 3.0 hours on LLaVA-v1.5-7B and 3.4 hours on LLaVA-v1.5-13B.
@@ 7.2 DPO Training Setups
7.3 Evaluation Benchmark Setups
We follow the original evaluation settings and benchmark splits for AMBER (Wang et al. 2023), MMHal (Sun et al. 2023), and OBJHal (Yu et al. 2024) as specified in their respective papers. For POPE (Li et al. 2023), we construct a new benchmark of 9,000 VQA pairs by sampling using the popular, random, and adversarial strategies.
For evaluation metrics, we adopt four response-level hallucination measures across different benchmarks: CHAIR (Rohrbach et al. 2018) for object hallucination detection, object coverage (Cover) for completeness measurement, response-level hallucination rate (Hal-Rate) for overall hallucination assessment, sentence-level hallucination rate (CRs) for holistic response evaluation, and object mention-level hallucination rate (CRi) for fine-grained object-level analysis.
For evaluation feedback collection, we employ the en-core-web-lg English NLP pipeline for AMBER to extract structured semantic cues as lightweight and reproducible evaluators. For MMHal and OBJHal, we utilize the expert GPT-4V model (Hurst et al. 2024) (gpt-4-1106-vision-preview) for feedback evaluation, following the established protocols of benchmarks.
8 Algorithm Flowchart
@@
9 Extended Ablation Studies
9.1 Impact of token-level perturbation magnitude
@@
9.2 Impact of frequency regularization
@@
[…]
@@
10 Additional Experimental Results
@@
11 Discussions and Insights
11.1 Why TARS outperforms DPO: Beyond numbers
@@
11.2 Token perturbation sensitivity
Impact of perturbation strength. TARS determines the number of perturbed tokens based on the model’s confidence and a scaling factor $\omega$. In Section 9.1, we conduct a systematic evaluation of perturbation magnitude to assess sensitivity. Results show that moderate perturbation levels (e.g., $\omega=1\mathrm{e}{-3}$) yield optimal hallucination suppression, while excessively small or large values lead to under- or over-regularization. These findings confirm that the model benefits from controlled perturbation, and that TARS is robust across a range of $\omega$ values when properly calibrated.
@@ 11.2 Token perturbation sensitivity
11.3 Spectrum-based alignment vs token-level alignment

### v3

@@ 4.1 Experiment Details
Intern-VL2.5-7B (Chen et al., 2024e)§ 7.9 54.7 37.1 3.2 3.54 0.26 - - 36.0 9.1
Qwen-VL2.5-8B (Bai et al., 2025)§ 4.6 54.6 21.1 1.3 3.29 0.27 - - 40.7 8.6
DeepSeek-VL2-27B (Wu et al., 2024)§ 2.4 56.6 16.3 0.9 2.84 0.27 - - 10.0 7.0
GPT-4o (Hurst et al., 2024)§ 2.5 60.9 17.6 0.8 3.87 0.24 - - 29.3 6.7
LLaVA-v1.5-7B (Liu et al., 2023b)§ 7.6 51.7 35.4 4.2 2.02 0.61 80.0 61.8 54.0 15.8
+ RLHF (Sun et al., 2024) 8.3 52.2 41.8 4.5 1.93 0.67 82.0 69.3 56.0 15.2
+ RLAIF (Yu et al., 2025) 3.0 50.3 16.5 1.0 2.89 0.42 88.1 88.0 13.7 4.2
+ HALVA (Sarkar et al., 2025a) 6.9 52.8 33.2 3.5 2.12 0.59 87.5 79.6 47.3 14.6
+ DPO (Li et al., 2024) 4.9 56.6 26.4 2.5 2.19 0.61 87.8 82.0 14.0 5.0
+ CHiP-DPO (Fu et al., 2025) 2.9 57.3 19.9 1.0 2.32 0.57 81.1 91.8 7.3 4.3
+ OPA-DPO (Yang et al., 2025) 2.7 47.4 12.5 0.9 2.78 0.46 87.4 86.2 13.3 4.5
@@
LLaVA-v1.5-13B (Liu et al., 2023b)§ 6.7 52.1 32.5 3.5 2.39 0.53 74.6 55.2 50.0 14.5
+ RLHF (Sun et al., 2024) 7.1 51.4 36.3 3.6 2.10 0.67 83.6 71.2 46.7 11.6
+ HALVA (Sarkar et al., 2025a) 6.5 53.4 30.1 3.3 2.28 0.56 86.8 75.6 42.7 12.1
+ DPO (Li et al., 2024) 4.1 56.7 24.3 2.2 2.48 0.50 85.2 84.3 19.0 7.2
+ CHiP-DPO (Fu et al., 2025) 3.8 58.6 20.8 1.7 2.70 0.46 86.6 74.9 30.0 6.2
+ OPA-DPO (Yang et al., 2025) 2.8 48.4 13.5 1.0 3.02 0.40 87.2 80.7 18.3 5.1
@@ 4.2 Evaluation on Hallucination Benchmarks
Table 2: Ablation results of token-level perturbation (TP), cross-modal alignment score (CAS), and spectral preference alignment (SPA) under 7B scale.
@@
Figure 5: Comparison of AMBER hallucination rate versus preference data scale.
@@ 4.3 Ablation Analyses on Component
(1) Token-level perturbation (TP) in Equation 6, which introduces distributional shifts and proves essential for revealing token-level vulnerabilities, its removal increases Cog from 0.4 to 2.5.
@@
LLaVA-v1.5 (Liu et al., 2023b) - -
+ RLHF (Sun et al., 2024) 122k self-reward
+ RLAIF (Yu et al., 2025) 16k LLaVA-Next
+ HALVA (Sarkar et al., 2025a) 22k GPT-4V
+ DPO (Li et al., 2024) 5k self-reward
+ CHiP-DPO (Fu et al., 2025) 5k self-reward
+ OPA-DPO (Yang et al., 2025) 4.8k GPT-4V
@@ 5 Related Work
Multimodal large language models (MLLMs) extend LLMs by integrating visual inputs to support multimodal reasoning (Chen et al., 2024b; Feng et al., 2025; Jain et al., 2024). Typically, visual features are extracted by a vision encoder, aligned through a connector, and processed by the LLM (Liu et al., 2023b; Parekh et al., 2024). Despite strong performance, MLLMs often produce factually incorrect or visually ungrounded outputs, undermining reliability (Bai et al., 2024; Chen et al., 2025). This issue is more severe than in unimodal LLMs (Chen et al., 2024d; Jiang et al., 2024a), mainly due to modality imbalance (Ma et al., 2024; He et al., 2024) and ineffective fusion (Bellagente et al., 2023; Ji et al., 2023). Recent studies attribute these failures to persistent misalignment between multimodal representations and human expectations, rather than model capacity (Chen et al., 2024c; Liu et al., 2024c; Ruan et al., 2025).
A key bottleneck in addressing MLLM hallucinations lies in aligning model outputs with human preferences for factual consistency. Unlike knowledge-intensive pretraining (Chang et al., 2024; McKinzie et al., 2024) and instruction tuning (Chen et al., 2024a; Liu et al., 2023b), recent methods typically leverage small-scale human preference data refined via reinforcement learning (Yu et al., 2024; Casper et al., 2023). Direct preference optimization (DPO)(Rafailov et al., 2023; Pi et al., 2024) has become a leading approach due to its simplicity and effectiveness, demonstrated in ChiP-DPO(Fu et al., 2025) and OPA-DPO (Yang et al., 2025). However, DPO’s reliance on limited data can cause overfitting to superficial linguistic cues (Setlur et al., 2024; Fu et al., 2025), leading to distributional rigidity and reduced adaptability to modality-specific semantics (Ouali et al., 2024; Song et al., 2024). These limitations call for more adaptive alignment strategies that capture token-level variability and cross-modal dependencies.
@@ 6 Conclusion
  • Section A details model configurations, training settings for DPO and TARS, and token perturbation procedures.
  • Section B presents the min-max optimization algorithm of TARS in pseudocode form.
  • Section C includes extended ablations analyzing the perturbation magnitude and spectral regularization strength.
  • Section D reports additional benchmark results and fine-grained hallucination metrics.
  • Section E discusses model behavior, including sensitivity and design insights.
  • Section F showcases qualitative comparisons on representative examples.
Appendix A Implementation Details
A.1 Base Model Setups
We evaluate our method on LLaVA-v1.5 (Liu et al., 2023b) models with 7B and 13B parameters. LLaVA-v1.5 adopts Vicuna-7B/13B (Chiang et al., 2023) as the language backbone and CLIP-ViT-L/14  (Radford et al., 2021) as the vision encoder. The vision encoder also serves as the similarity function $\mathcal{G}(\cdot)$ used in Eq. (9) to compute alignment between visual inputs and text tokens. All experiments are conducted using greedy decoding with a temperature of 0 to ensure deterministic outputs and reproducibility.
A.2 DPO Training Setups
[…]
@@ A.2 DPO Training Setups
A.3 Evaluation Benchmark Setups
We follow the original evaluation settings and benchmark splits for AMBER (Wang et al., 2023), MMHal (Sun et al., 2023), and OBJHal (Yu et al., 2024) as specified in their respective papers. For POPE (Li et al., 2023), we construct a new benchmark of 9,000 VQA pairs by sampling using the popular, random, and adversarial strategies.
For evaluation metrics, we adopt four response-level hallucination measures across different benchmarks: CHAIR (Rohrbach et al., 2018) for object hallucination detection, object coverage (Cover) for completeness measurement, response-level hallucination rate (Hal-Rate) for overall hallucination assessment, sentence-level hallucination rate (CRs) for holistic response evaluation, and object mention-level hallucination rate (CRi) for fine-grained object-level analysis.
For evaluation feedback collection, we employ the en-core-web-lg English NLP pipeline for AMBER to extract structured semantic cues as lightweight and reproducible evaluators. For MMHal and OBJHal, we utilize the expert GPT-4V model (Hurst et al., 2024) (gpt-4-1106-vision-preview) for feedback evaluation, following the established protocols of benchmarks.
Appendix B Algorithm Flowchart
@@
Appendix C Extended Ablation Studies
C.1 Impact of token-level perturbation magnitude
@@
C.2 Impact of frequency regularization
@@
We evaluate $\lambda\in\{0.01,0.02,0.05,0.10,0.20,0.50,1.00\}$ and present results in Table 5. We omit $\lambda=0.00$ here as it reduces the method to DPO with perturbation. As shown in the table, introducing spectral alignment leads to consistent improvements across all benchmarks. Performance improves steadily as $\lambda$ increases from 0.01 to 0.20, with hallucination rates (CHAIR, Hal-Rate, $_{\text{s}}, CR$$_{\text{i}}$) decreasing and coverage improving. The best trade-off is achieved at $\lambda=0.20$, where TARS achieves the lowest hallucination and strongest grounding.
@@
Appendix D Additional Experimental Results
@@
Appendix E Discussions and Insights
E.1 Why TARS outperforms DPO: Beyond numbers
@@
E.2 Token perturbation sensitivity
Impact of perturbation strength. TARS determines the number of perturbed tokens based on the model’s confidence and a scaling factor $\omega$. In Section C.1, we conduct a systematic evaluation of perturbation magnitude to assess sensitivity. Results show that moderate perturbation levels (e.g., $\omega=1\mathrm{e}{-3}$) yield optimal hallucination suppression, while excessively small or large values lead to under- or over-regularization. These findings confirm that the model benefits from controlled perturbation, and that TARS is robust across a range of $\omega$ values when properly calibrated.
@@ E.2 Token perturbation sensitivity
E.3 Spectrum-based alignment vs token-level alignment